# Precision Mars Entry Navigation with Atmospheric Density Adaptation via Neural Networks

### v1

@@ Abstract
Discrepancies between the true Martian atmospheric density and the onboard density model can significantly impair the performance of spacecraft entry navigation filters. This work introduces a new approach to online filtering for Martian entry by using a neural network to estimate atmospheric density and employing a consider analysis to account for the uncertainty in the estimate. The network is trained on an exponential atmospheric density model, and its parameters are dynamically adapted in real time to account for any mismatches between the true and estimated densities. The adaptation of the network is formulated as a maximum likelihood problem, leveraging the measurement innovations of the filter to identify optimal network parameters. The incorporation of a neural network enables the use of stochastic optimizers known for their efficiency in the machine learning domain within the context of the maximum likelihood approach. Performance comparisons against previous approaches are conducted in various realistic Mars entry navigation scenarios, resulting in superior estimation accuracy and precise alignment of the estimated density with a broad selection of realistic Martian atmospheres sampled from perturbed Mars-GRAM data.
@@ 1 Introduction
Among the most demanding challenges encountered in spacecraft navigation is atmospheric entry, particularly Martian entry. The entry phase is characterized by its intense dynamics, scarcity of available measurements, and uncertain atmospheric information [1]. While knowledge of the Martian atmosphere, weather, and seasons has improved over the last decades, it is not as accurately predicted as the atmosphere of Earth. Large uncertainties in models used to estimate the Martian atmosphere can make navigation during the entry phase challenging and can significantly degrade guidance solutions affecting landing accuracy [2]. The development of standard Entry, Descent, and Landing (EDL) systems is additionally complicated by atmospheric variability across Martian years. These systems typically employ a correction mechanism, adjusting a nominal exponential profile with a multiplicative factor based on the drag force experienced by the spacecraft [3]. However, constraining density estimation to an exponential profile poses challenges due to the significant atmospheric dust content on Mars, known to increase the temperature of the lower atmosphere and subsequently reduce its density [4]. As a result of this, future human and robotic Mars exploration missions require an improvement in navigation solutions by several orders of magnitude over the current state-of-the-art [5].
@@
Traditionally, Mars entry missions have used inertial measurement units (IMU) for navigation, in lieu of onboard models of non-conservative forces. However, the drawback of IMU-based solutions lies in their tendency to drift over time, especially in the absence of external measurements. When employed for state correction rather than propagation, IMU measurements alone can result in rendering most estimation states unobservable [6]. Due to this, different techniques have been explored to improve the navigation solutions during Martian entry. In early efforts of using only IMU measurements for entry navigation, Zanetti and Bishop [7] successfully quantified the accuracy of state estimates through dead-reckoning during a standard Mars atmospheric entry. Following this approach, Lou et al. [8] presented a strategy attempting to mitigate the effects of unobservable uncertain parameters through consider analysis during entry, achieving promising results. Additionally, Jiang et al. [9] introduced a more accurate multi-sensor strategy, where instead of only using an IMU for navigation, information from real-time flush air data system, a single orbiting radio beacon, and an IMU were fused using an unscented Kalman filter, providing better estimation results. More recently, the integration of aerodynamic pressure and aerothermal sensor data for real-time navigation has been proposed [10]. Nevertheless, most of these solutions encounter a common challenge: the use of an onboard exponential atmospheric density model, or reliance on external measurements like orbiting or surface beacons. These assumptions may pose difficulties when faced with real density profiles, potentially leading to filter divergence. Furthermore, it is worth noting that due to the presence of the heat-shield cover, onboard optical or radio navigation may not be available during entry [10]. Currently, a promising research area for entry navigation involves addressing the atmospheric uncertainty, offering potential improvements in estimation accuracy for Mars entry missions.
@@
The descent and landing phases are typically supported by sensors such as altimeters, velocimeters, or LIDARs, which can aid the navigation solution [11, 12]. Since these sensors are not available during the entry phase, this segment presents the most significant challenge for navigation performance. Due to the limited availability of sensors during entry, accurate density estimation becomes crucial. Correctly estimating atmospheric density and quantifying its inherent uncertainty can greatly improve navigation outputs, thus enhancing overall precision landing performance. A range of methodologies have been investigated to address the atmospheric uncertainties within the Martian environment, with the aim of improving the performance of EDL systems. These approaches can be categorized into two main groups: offline reconstructions, which involves working with real data acquired during the entry phase, and online estimation techniques designed for use during the entry phase. Dutta et al. [13] conducted an analysis of various estimation techniques, including the extended Kalman filter (EKF) and the unscented Kalman filter (UKF) for offline reconstruction. In their work, the main objective was to reconstruct the vehicle’s trajectory, aerodynamic coefficients, and the atmospheric profile experienced by the system in an offline context. Subsequently, in their follow-up work [14], they demonstrated how adaptive filtering techniques, such as covariance matching [15], could enhance the estimation performance of aerodynamic coefficients and atmospheric properties. These advanced techniques were later employed to successfully reconstruct atmospheric properties using real data acquired by the Mars Science Laboratory (MSL) [16].
@@
To improve the precision of filtering techniques in online estimation for EDL, various strategies have also been explored. A common solution involves carrying a nominal density onboard the filter and augmenting the system state to estimate a correction factor that accounts for the deviation between the nominal density and the actual density experienced by the system [10, 2]. In addition, machine learning approaches have also been used for density estimation in EDL problems. Wagner et al. [17] used an ensemble of neural networks to dynamically adapt to real-time changes in atmospheric properties during Martian entry. Additionally, Amato and McMahon [18] employed a long short-term memory network to estimate atmospheric density based on inertial measurements and guidance commands. Most recently, Roelke et al. [19] introduced two density profile prediction techniques for aerocapture, using density histories computed from navigation measurements to enhance targeting accuracy.
@@
In this work, we propose a new approach to onboard filtering for Martian entry by using a neural network (trained pre-flight on a simplified model retrieved from Mars-GRAM trajectories) to estimate atmospheric density and accounting for atmospheric uncertainty through a consider approach. The atmospheric mismatch between the true density and the model used by the filter is addressed by adapting the parameters of the network in real-time using the measurement innovations of the navigation filter. The adaptation of the network is performed using aerodynamic pressure measurements, aerothermal sensor data, and an IMU, mirroring sensor configurations from prior Martian EDL missions [20, 21, 22]. By leveraging the capabilities of state-of-the-art stochastic optimizers, nonlinear activation functions, and automatic differentiation, the adaptation step is posed as a maximum likelihood (ML) problem, resulting in consistent filtering performance and accurate estimation of the true density.
@@
The remainder of this paper is organized as follows: First, a background of Martian entry is provided in Section 2. Previous approaches to density adaptation are introduced in Section 3, and the approach to density adaptation using neural networks is presented in Section 4. The performance of the new approach is compared to previous approaches on a realistic Mars entry navigation scenario in Section 5. Finally, Section 6 provides conclusions and outlines future work opportunities.
@@ 2 Martian Entry
This section provides insight into the dynamics, measurement, and atmospheric density models used in Martian entry. The aim of this analysis is to emphasize the importance of acquiring an accurate atmospheric density estimate during the entry phase. A clear understanding of these fundamentals is important for recognizing the crucial role that accurate density estimation plays in mission success [1].
@@ 2.1 Dynamics
The dynamics of entry vehicles during Martian entry are often represented using a simplified three degrees-of-freedom (3-DOF) model. In this approach and for the entirety of this work, point mass dynamics are assumed, meaning that no consideration is given to attitude dynamics - a common convention observed in prior literature [23, 6, 24]. The position of the vehicle is described in terms of planet-centric coordinates, which include the planet-centric radius (), latitude (), and longitude () [24],
The use of planet-centric latitude and longitude ( and ) facilitates the application of spherical equations of motion, even when dealing with a non-spherical planet, as the radius of the planet is allowed to vary with latitude and longitude [24]. The velocity (), flight path angle (), and heading angle () are defined relative to the planet surface and are based on the vehicle-carried local horizontal frame (LH). The flight path angle is defined as negative-down and the heading angle is defined in the horizontal plane where is due East and due North is [24]. The equations describing the time evolution of the velocity-related terms are given by:
where represents the planetary rotation rate, is the planetary gravity, is the bank angle, and the lift () and drag () accelerations are defined as follows:
@@
Given the substantial hypersonic speeds associated with the entry phase, wind velocities are considerably lower compared to the velocity of the vehicle, i.e., [18]. In such cases, it is reasonable to make the approximation that . Moreover, for brief-duration flights such as the entry phase in EDL, the impact of planetary rotation can be deemed negligible [10]. Hence, by omitting terms related to planetary rotation and assuming a first-order gravity model, the equations governing velocity parameters simplify to:
@@ 2.2 Onboard Sensors
The measurement model of the onboard sensors typically used in EDL missions is closely tied to the atmospheric density estimate as well. In this work, three distinct onboard sensors are under consideration. These sensors encompass an IMU, a cluster of pressure sensors strategically placed on the heat shield of the vehicle, and a suite of thermocouples situated on the forebody of the vehicle. These sensors are carefully selected to replicate the sensor configurations used in past Martian EDL missions [10, 20, 21, 22]. The IMU measures the non-gravitational accelerations encountered by the entry vehicle. In this work, these measurements are conducted relative to the body frame, as follows:
where represents measurement noise, is the transformation matrix to change coordinates from velocity frame to body frame, and are the non-gravitational accelerations in the velocity frame defined as:
The pressure sensors are assumed to offer a reasonably accurate evaluation of the stagnation-point pressure, while the thermocouples are assumed to provide real-time measurements of the heating rate. It is important to acknowledge that the precision of these measurements may vary according to the flight regime. For instance, pressure transducers optimized for hypersonic flight may not deliver satisfactory results during supersonic flight. Nevertheless, this work assumes the availability of measurements throughout the entire entry stage. The dynamic pressure measurement, assuming an aggregate measurement from all relevant sensor arrays, is expressed as follows [10]:
The measured convective heating rate is modeled by the Sutton-Graves relation, such that [25, 26]:
where kgm, is the vehicle nose radius and is measurement noise. These equations further emphasize the critical importance of having an accurate atmospheric density estimate, as they all rely on this fundamental parameter. This connection illustrates how easily a navigation system can diverge when confronted with inaccurate estimates or inaccurate density models stored in the onboard computer of the vehicle, emphasizing the critical role of obtaining a highly accurate density estimate for mission success and safety.
@@ 2.3 Density Models
To model Martian atmospheric density, various approaches are available, ranging from more intricate models for increased accuracy to simpler equations for faster estimates, although with reduced precision. The most straightforward equation for modeling atmospheric density is based on an exponential law, assuming isothermal equilibrium [27]. This model is represented by the equation:
@@ 3 Approaches to Density Adaptation
As models cannot provide a perfect estimation of the atmospheric density, various adaptive techniques can be used to address potential errors or unforeseen behavior that the vehicle may encounter during the entry phase. The first approach involves covariance matching, wherein adjustments are made to the process noise and measurement covariance matrices to align with the state and measurement innovations in the navigation filter. The second technique entails the estimation of correction parameters, used to rectify a nominal density profile established during the offline phases of the navigation filter design. Lastly, multiple model adaptive estimation can also be used for this purpose, where a bank of estimators operates in parallel, yielding a more robust estimate by considering different density profiles. This section delves into the presentation and discussion of each of these approaches.
Covariance matching techniques have found utility in atmospheric reconstruction due to the inherent uncertainty in atmospheric parameters, where the process noise and measurement covariances matrices are often unknown. When directly estimating the density, that is, the density is part of the estimation state, covariance matching techniques have been shown to be advantageous [14]. It is worth noting that incorporating the density as part of the state introduces constraints, notably on the dynamics, as the equation for the rate of change is derived from the hydrostatic equation and the perfect gas law [13].
@@
Another approach to atmospheric density adaptation includes the correction to a nominal profile by augmenting the estimation state [2]. In this solution, a nominal density profile () is determined offline and carried by the onboard navigation system. The nominal profile is corrected in an online fashion to account for potential errors. To correct the nominal profile, different approaches can be used. For example, the estimated state can be augmented with a correction factor for the density [2],
@@
With this approach, the estimated density can be calculated as:
Estimating in this manner behaves more favorably from a numerical perspective because the estimator primarily seeks to refine an existing approximate density profile rather than directly estimating the density [2]. The estimated value will tend to hover near one, regardless of altitude, when the nominal density profile closely approximates the actual density experienced by the vehicle. A different way of correcting a nominal profile involves directly modeling the atmosphere as in (16) and estimating the density at the reference radial position () and the atmospheric scale height () [10], thus:
However, this solution can be ill-conditioning since the exponential model can amplify errors at higher altitudes [14]. It is important to note that augmentation approaches perform the adaptation to the atmospheric density as a linear function of the measurement, since the extended Kalman filter (EKF) and the unscented Kalman filter (UKF) are typically used in these scenarios.

### v2

@@ Abstract
Spacecraft entering Mars require precise navigation algorithms capable of accurately estimating the vehicle’s position and velocity in dynamic and uncertain atmospheric environments. Discrepancies between the true Martian atmospheric density and the onboard density model can significantly impair the performance of spacecraft entry navigation filters. This work introduces a new approach to online filtering for Martian entry using a neural network to estimate atmospheric density and employing a consider analysis to account for the uncertainty in the estimate. The network is trained on an exponential atmospheric density model, and its parameters are dynamically adapted in real time to account for any mismatch between the true and estimated densities. The adaptation of the network is formulated as a maximum likelihood problem by leveraging the measurement innovations of the filter to identify optimal network parameters. Within the context of the maximum likelihood approach, incorporating a neural network enables the use of stochastic optimizers known for their efficiency in the machine learning domain. Performance comparisons are conducted against two online adaptive approaches, covariance matching and state augmentation and correction, in various realistic Martian entry navigation scenarios. The results show superior estimation accuracy compared to other approaches, and precise alignment of the estimated density with a broad selection of realistic Martian atmospheres sampled from perturbed Mars-GRAM data.
@@ 1 Introduction
Among the most demanding challenges encountered in spacecraft navigation is atmospheric entry on Mars. The entry phase is characterized by intense dynamics, a scarcity of available measurements, and uncertain atmospheric information [1]. Due to the uncertainty in atmospheric conditions, discrepancies between the true atmospheric density and onboard models can make navigation during entry challenging and significantly degrade guidance solutions, affecting landing accuracy [2]. Despite improvements in the knowledge of the Martian atmosphere, its high variability, coupled with a scarcity of comprehensive data, makes it less accurately predictable compared to Earth’s atmosphere [1], complicating the development of standard Entry, Descent, and Landing (EDL) systems. Standard EDL systems typically use a correction mechanism that adjusts an exponential density profile with a multiplicative factor based on the drag force experienced by the spacecraft [3]. However, constraining density estimation to an exponential profile can pose challenges for Martian entry. For example, the significant atmospheric dust content on Mars is known to increase the temperature of the lower atmosphere and subsequently reduce its density [4]. As a result, future human and robotic Mars exploration missions require improvements by several orders of magnitude over the current state-of-the-art [5].
@@
Previous Mars entry missions have used inertial measurement units (IMU) for navigation instead of onboard models of non-conservative forces. However, IMU-based solutions tend to drift over time, especially without external measurements. If standalone IMU measurements are used for state correction rather than propagation, most estimation states can become unobservable during entry [6]. Therefore, various techniques have been explored to improve navigation solutions during Martian entry. Zanetti and Bishop [7] successfully quantified the accuracy of state estimates through dead-reckoning during standard Martian atmospheric entry only using IMU measurements for navigation. Following this approach, Lou et al. [8] presented a strategy attempting to mitigate the effects of unobservable uncertain parameters through consider analysis during entry. Additionally, Jiang et al. [9] introduced a more accurate multisensor strategy, where instead of using only an IMU for navigation, information from a real-time flush air data system, a single orbiting radio beacon, and an IMU were fused using an unscented Kalman filter, providing better estimation results. Recently, the integration of aerodynamic pressure and aerothermal sensor data for real-time navigation has been proposed [10]. However, most of these solutions face a common challenge: they either rely on an onboard exponential atmospheric density model or depend on external measurements, such as orbiting or surface beacons. The assumption of an exponential density model may pose difficulties when faced with real density profiles, potentially leading to filter divergence. Furthermore, due to the presence of a heat-shield cover on the entry vehicle, onboard optical or radio navigation may not be available during entry [10], making external measurements unavailable.
@@
The descent and landing phases are typically supported by sensors, such as altimeters, velocimeters, or LIDARs, which can aid navigation solutions [11, 12]. However, these sensors are not available during the entry phase, which poses the most significant challenge for navigation. Accurate density estimation becomes crucial in the absence of additional sensors. Therefore, effectively estimating atmospheric density and quantifying its uncertainty can significantly enhance navigation outputs, improving precision landing performance. A range of methodologies have been investigated to address atmospheric uncertainties within the Martian environment with the aim of improving the performance of EDL systems. These approaches can be categorized into two main groups: offline reconstructions, which involve working with real data acquired during the entry phase, and online estimation techniques, which are designed for use during the entry phase. Dutta et al. [13] analyzed various estimation techniques, including the extended Kalman filter (EKF) and unscented Kalman filter (UKF) for offline reconstruction. In their work, the main objective was to reconstruct the vehicle trajectory, aerodynamic coefficients, and atmospheric profile experienced by the system in an offline context. Subsequently, in their follow-up work [14], they demonstrated how adaptive filtering techniques such as covariance matching [15, 16] could enhance the estimation performance of aerodynamic coefficients and atmospheric properties. These advanced techniques were later employed to successfully reconstruct atmospheric properties using real data acquired by the Mars Science Laboratory (MSL) [17].
@@
To improve filtering precision in online estimation for EDL, various strategies have been explored. A common solution involves carrying a nominal density onboard the filter and augmenting the system state to estimate a correction factor that accounts for the deviation between the nominal density and actual density experienced by the system [10, 2]. In addition, machine learning approaches have been used for density estimation in EDL problems. Wagner et al. [18] used an ensemble of neural networks to dynamically adapt to real-time changes in atmospheric properties during Martian entry. Additionally, Amato and McMahon [19] employed a long short-term memory network to estimate atmospheric density based on inertial measurements and guidance commands. Most recently, Roelke et al. [20] introduced two density profile prediction techniques for aerocapture, using density histories computed from navigation measurements to enhance targeting accuracy.
@@
In this work, we propose a new approach to onboard adaptive filtering for Martian entry by using a neural network, trained pre-flight on a simplified model retrieved from Mars-GRAM trajectories, to estimate atmospheric density. To address the atmospheric mismatch between the true density and the neural network, the network’s parameters are adapted in real time using measurement innovations. The measurement innovations are defined as the difference between the true measurements and the predicted measurements of the navigation filter. Furthermore, we account for the uncertainty in the atmospheric density estimate using consider analysis. Consider analysis acknowledges errors in the dynamic and measurement models due to uncertain parameters [21], such as atmospheric density in this case. This approach enables the navigation filter to correctly weight the data while accounting for the correlation between the state and the uncertain parameter [22]. The adaptation of the network is performed using aerodynamic pressure measurements, aerothermal sensor data, and an IMU, mirroring sensor configurations from prior Martian EDL missions [23, 24, 25]. Leveraging the capabilities of state-of-the-art stochastic optimizers, nonlinear activation functions, and automatic differentiation, the adaptation step is posed as a maximum likelihood (ML) problem, resulting in consistent filtering performance and accurate estimation of the true density.
@@
The remainder of this paper is organized as follows: First, a background of Martian entry is provided in Section 2. Previous approaches to density adaptation are introduced in Section 3, and the approach to density adaptation using neural networks is presented in Section 4. The performance of the new approach is compared to previous approaches on a realistic Mars entry navigation scenario in Section 5. Section 6 discusses some of the limitations of the framework presented and outlines future work opportunities. Finally, Section 7 presents the conclusions of the work.
@@ 2 Martian Entry
This section provides insight into the dynamics, measurement, and atmospheric density models used in Martian entry. The aim of this analysis is to emphasize the importance of acquiring an accurate atmospheric density estimate during the entry phase.
@@ 2.1 Dynamics
The dynamics of entry vehicles during Martian entry are often represented using a simplified three degrees-of-freedom (3-DOF) model. In this approach and for the entirety of this work, point mass dynamics are assumed, meaning that no consideration is given to attitude dynamics, a common convention observed in prior literature [26, 6, 27, 28, 10]. The position of the vehicle is described in terms of planet-centric radius (), latitude (), and longitude () which are shown with respect to the Mars-centered Mars-fixed (MCMF) frame [6] in Fig. 1. The MCMF frame is centered and fixed to the planet aligning with the axis of rotation. These coordinates allow the use of spherical equations of motion, even for non-spherical planets, as the radius can vary with latitude and longitude [27]. The dynamic equations used to model the evolution of these variables with respect to time are given by [6, 28, 10],
The magnitude of the velocity vector (), flight path angle (), and heading azimuth () are used to describe the velocity of the vehicle as shown with respect to the Geographic frame [6] in Fig. 1. The Geographic frame is centered on the vehicle and , , are aligned with Mars geographic directions North, East and Nadir, respectively. The flight path angle is defined as negative-down, and the heading azimuth is defined in the horizontal plane with as due North and as due East. The equations describing the time evolution of the velocity-related terms are given by [6, 28, 10]:
where represents the planetary rotation rate, is the planetary gravity, is the bank angle, and the lift () and drag () accelerations are defined as follows [6]:
@@
Given the substantial hypersonic speeds associated with the entry phase, wind velocities are considerably lower compared to the velocity of the vehicle, i.e., [19]. In such cases, it is reasonable to make the approximation that . Moreover, for brief-duration flights such as the entry phase in EDL, the impact of planetary rotation can be deemed negligible [10]. Note that neglecting planetary rotation makes frame a simple inertial frame. Hence, by omitting terms related to planetary rotation and assuming a first-order gravity model, the equations governing velocity parameters simplify to [29]:
@@ 2.2 Onboard Sensors
The measurement model of the onboard sensors typically used in EDL missions is closely tied to the atmospheric density estimate as well. In this work, three distinct onboard sensors are under consideration. These sensors encompass an IMU, a cluster of pressure sensors strategically placed on the heat shield of the vehicle, and a suite of thermocouples situated on the forebody of the vehicle. These sensors are carefully selected to replicate the sensor configurations used in past Martian EDL missions [10, 23, 24, 25].
2.2.1 Inertial measurement unit
The IMU measures the non-gravitational accelerations encountered by the entry vehicle. In this work, these measurements are conducted relative to the body frame aligned with the axes of symmetry in the vehicle [10]. The acceleration measurements are described by:
where represents measurement noise, is the transformation matrix to change coordinates from velocity frame to body frame, and are the non-gravitational accelerations in the velocity frame defined as [10]:
The velocity frame is centered on the vehicle, with its -axis aligned in the direction of the flight path angle, its -axis aligned with the local vertical plane orthogonal to the -axis, and the -axis completing the right-handed system [10].
2.2.2 Cluster of pressure sensors
The pressure sensors are assumed to offer a reasonably accurate evaluation of the stagnation-point pressure. It is important to acknowledge that the precision of these measurements may vary according to the flight regime. For instance, pressure transducers optimized for hypersonic flight may not deliver satisfactory results during supersonic flight. Nevertheless, this work assumes the availability of measurements throughout the entire entry stage. The dynamic pressure measurement, assuming an aggregate measurement from all relevant sensor arrays, is expressed as follows [10]:
where represents measurement noise.
2.2.3 Suite of thermocouples
The suite of thermocouples is assumed to provide real-time measurements of the convective heating rate. The measured convective heating rate, assuming again an aggregate measurement from all relevant sensor arrays, is modeled by the Sutton-Graves relation [30, 31],
where kg1/2m-1, is the vehicle nose radius and is measurement noise.
These equations further emphasize the critical importance of having an accurate atmospheric density estimate, as they all rely on this fundamental parameter. This connection illustrates how easily a navigation system can diverge when confronted with inaccurate estimates or inaccurate density models stored in the onboard computer of the vehicle, emphasizing the role of obtaining a highly accurate density estimate for mission success and safety.
@@ 2.3 Density Models
To model Martian atmospheric density, various approaches are available, ranging from more intricate models for increased accuracy to simpler equations for faster estimates, although with reduced precision.
2.3.1 Exponential models
The most straightforward equation for modeling atmospheric density is based on an exponential law, assuming isothermal equilibrium [32]. This model is represented by the equation:
@@ 3 Approaches to Density Adaptation
As models cannot provide a perfect estimation of the atmospheric density, various adaptive techniques can be used to address potential errors or unforeseen behavior that the vehicle may encounter during the entry phase. The first approach involves covariance matching, wherein adjustments are made to the process and measurement noise covariance matrices to align with the state and measurement innovations in the navigation filter. The second technique entails the estimation of correction parameters, used to rectify a nominal density profile established during the offline phases of the navigation filter design. Lastly, multiple model adaptive estimation can also be used for this purpose, where a bank of estimators operates in parallel, yielding a more robust estimate by considering different density profiles. This section delves into the presentation and discussion of each of these approaches.
3.1 Covariance Matching
Covariance matching techniques have found utility in atmospheric reconstruction due to the inherent uncertainty in atmospheric parameters, where the process and measurement noise covariances matrices are often unknown. When directly estimating the density, that is, the density is part of the estimation state, covariance matching techniques have been shown to be advantageous [14]. It is worth noting that incorporating the density as part of the state introduces constraints, notably on the dynamics, as the equation for the rate of change is derived from the hydrostatic equation and the perfect gas law [13].
@@ 3.1 Covariance Matching
3.2 State Augmentation and Correction
Another approach to atmospheric density adaptation includes the correction to a nominal profile by augmenting the estimation state [2, 10]. In this solution, a nominal density profile () is determined offline and carried by the onboard navigation system. The nominal profile is corrected in an online fashion to account for potential errors. To correct the nominal profile, different approaches can be used. For example, the estimated state can be augmented with a correction factor for the density [2],
@@ 3.2 State Augmentation and Correction
With this approach, the correction factor can be estimated within the navigation filter and the estimated density can be calculated as:
Estimating behaves favorably from a numerical perspective, because the estimator primarily seeks to refine an existing approximate density profile rather than directly estimating the density [2]. The estimated value tends to hover near one, regardless of altitude, when the nominal density profile closely approximates the actual density experienced by the vehicle. Previous guidance approaches have estimated this correction factor by using the sensed deceleration measurements of the vehicle and applying a low-pass ad hoc filter [20]. When a Kalman filter is used for navigation, augmenting the filtering state with this correction factor effectively takes care of applying a low-pass filter.
A different way of correcting a nominal profile involves directly modeling the atmosphere as in (16) and estimating the density at the reference radial position () [29, 10] and the atmospheric scale height (), thus,
However, constraining the nominal density model to an exponential form can be ill-conditioned because the exponential model tends to amplify errors at higher altitudes [14]. It is important to note that augmentation approaches perform the adaptation to the atmospheric density as a linear function of the measurement, since the EKF and the UKF are typically used in these scenarios.
3.3 Others